OpenStudy (anonymous):

Solve 1 + 5/(x-1) >or equal to 7/6

OpenStudy (anonymous):

\[1+\frac{5}{x-1}\geq \frac{7}{6}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

subtract 1 to get \[\frac{5}{x-1}\geq \frac{1}{6}\] as a first step

OpenStudy (anonymous):

the flip it and change the inequality \[\frac{x-1}{5}\leq 6\]

OpenStudy (anonymous):

x< or equal to 31 thank you!

OpenStudy (anonymous):

last two steps a) multiply by 5 b) add 1 \[\frac{x-1}{5}\leq 6\\ x-1\leq 30\\ x\leq 31\]

OpenStudy (anonymous):

yeah you got it yw